OpenStudy (baldymcgee6):

Writing limits in latex, sometimes the x->c is under the "lim", but sometimes it is to the side of the "lim".

OpenStudy (baldymcgee6):

\[\large{\color{red}{\lim_{x\to 0^+} \left[(\tan (x))(\ln (x))\right]}}\]

hartnn (hartnn):

\(\large{\color{red}{\lim_{x\to 0^+} \left[(\tan (x))(\ln (x))\right]}}\)

hartnn (hartnn):

:(

hartnn (hartnn):

i just copy-pasted

OpenStudy (baldymcgee6):

why does it do that?!

OpenStudy (baldymcgee6):

in my assignment program i have the exact same code and in one it is under it, and the other it is to the side.,..

hartnn (hartnn):

no idea

OpenStudy (baldymcgee6):

hmmmm, odd

hartnn (hartnn):

when I write, i ALWAYS get \(\large{\color{red}{\lim_{x\to 0^+} \left[(\tan (x))(\ln (x))\right]}}\)

OpenStudy (baldymcgee6):

i figured it out... You have to be in display style.

hartnn (hartnn):

\[\(\large{\color{red}{\lim_{x\to 0^+} \left[(\tan (x))(\ln (x))\right]}}\)\]

hartnn (hartnn):

\[\large{\color{red}{\lim_{x\to 0^+} \left[(\tan (x))(\ln (x))\right]}}\]

hartnn (hartnn):

maybe i figured at the very same moment u did!

OpenStudy (baldymcgee6):

:D thanks for the help again!

hartnn (hartnn):

i didn't use displaystyle

hartnn (hartnn):

in first few cases i used `\(....\)` in last comment i used `\[....\]`

OpenStudy (baldymcgee6):

okay! cool, thanks!